### v2

# Analysis Theory of Data Economy : Dataization, Technological Progress and Dynamic General Equilibrium

Abstract
This paper constructs a clean and efficient representative agent model of the data economy from a macroeconomics perspective, in order to analyze the impact of dataization and technological progress on the dynamic general equilibrium of “consumption-capital”, and the catalysis effect of dataization on technological progress. We first set the data in production comes from dataization of the total output of the society and is exponentially functionally related to the technology. Secondly, the data production function is used to solve the optimization problem for firms and households and to construct a dynamic general equilibrium of “consumption-capital” based on the endogenous interest rate solved by maximizing the returns of firms. Finally, by using numerical simulation and phase diagram analysis, we find that the effects of increasing dataization and encouraging technological progress each exhibit different nonlinear characteristics for equilibrium capital and equilibrium consumption, we thus conclude that dataization enables the positive effects of technological progress on economic development to be more rapid and persistent. We select two types of Chinese policies regarding data openness to represent the role of dataization, and demonstrate the catalysis role of dataization for the development of the digital economy by setting up difference in difference (DID) experiments, which provide persuasive evidence for the theoretical interpretation of the paper.
Key Words: Data economy, Dataization, Technological progress, Representative agent model, Dynamic general equilibrium
JEL Codes: C02, C80, D11, D21, E2, O11
1 Introduction
Data essentially shifts the material basis of knowledge and information production from traditional discourse practices to quantifiable and tradable data resources, completing the dataization transformation of knowledge, information and production factors. The special economic attributes of data as a factor of production, such as non-rivalry, replicability, virtuality, open source, and non-depletion at the physical level (Jones and Tonetti, 2020; Goldfarb and Tucker, 2019; Lynch, 2008), enable it to be integrated into the cycle of capital accumulation in a way of decreasing marginal cost and increasing returns to scale.
Much of the current research on the impact of data on economic systems focuses on how data drive productivity leaps and economic growth, exploring the role of data elements empowering digital-intelligent technologies, in terms of productivity enhancement, especially total factor productivity, and the reduction effect on various production costs (Nose, 2023; Acemoglu and Restrepo, 2018; Graetz and Michaels, 2018), as well as treating data as intangible capital and constructing economic models containing data capital to assess the impact of big data on productivity (Corrado et al., 2024). Moreover, data elements can effectively drive the innovation of digital products (Athey, 2025), and not only that, since researching more novel problems can contribute to the evolution of knowledge, thus triggering the research cycle (Carnehl and Schneider, 2025), AI technologies empowered by data elements can significantly increase the amount and accuracy of scientific knowledge acquisition in a given domain, as a result, the desire of researchers to innovate and the efficiency of technological innovation can be effectively improved, and research on more novel problems can be facilitated (Gans, 2025a; Costa and Aparicio, 2025; Costa et al., 2024).At the same time, the use of data elements can also play a role in driving high-quality transformation of industries and optimizing the allocation of production resources (Acemoglu, 2025; Glaeser et al., 2018; Zhang et al., 2018; Schaefer et al., 2014; Einav and Levin, 2014).
There have also been studies of data-driven productivity based on the microeconomic attributes of the data factors : as a result of the revolution in production by digital technology, data factors are transformed more into digitized knowledge, which improves the production chain, regulates the rationing of factors of production, drives technological innovations, and raises the marginal productivity of the traditional factors of production, which in turn raises total output (Abis and Veldkamp, 2024). Through its own “multiplier effect”, data empowers traditional factors of production and can contribute to capital- and labor-expanding technological advances individually, or simultaneously, thereby driving increases in the marginal productivity of capital and labor factors, and thus increasing total factor productivity (Hémous and Olsen, 2022; Bessen, 2019; Acemoglu and Restrepo, 2018). Viewing data itself as a special kind of labor can help alleviate problems in terms of unequal returns to the data economy and restore a normal market for user contributions (Arrieta-Ibarra et al., 2018). However, the use of a “data” workforce is not always beneficial, and when firms overuse data-driven automation to eliminate traditional human capital, this leads to borrowing and redistributive frictions, which affects the efficiency of society’s operations (Beraja and Zorzi, 2025). Moreover, Brynjolfsson et al. (2025), from a heterogeneous worker perspective, find that data-driven AI technology can improve both the productivity and quality of work for low-qualified workers, but for high-qualified workers, it only slightly improves the productivity of their work, while the quality of their work decreases. Aum and Shin (2025) show that big data-enabled digital technologies have a significant negative impact on high-skilled workers, and while this can displace part of the labor force at the task performance level, it can also drive labor demand towards newer skill sets, thereby boosting society’s total labor demand (Hampole et al., 2025). However, research by Jiang et al. (2025) suggests that the relationship between data-driven AI technologies and the human workforce is complementary rather than substitutive, and that while this will lead to a significant increase in society’s marginal productivity and serve as an effective proxy for leadership, decision-making, and teamwork (Weidmann et al., 2025), it also disrupts the work-life balance of individuals. In addition, Farboodi et al. (2019) and Farboodi and Veldkamp (2021) regard data as an information resource that can reduce production uncertainty, and firms use data elements to acquire forward-looking knowledge, thus improving their prediction accuracy of the optimal production technology, which in turn enhances the level of productivity. The findings of Comin and Mestieri (2018) and Varian (2018) indicate that the diffusion effect of data elements on an economy is similar to a Logistic function, which is determined by the rate of diffusion of data elements in the economy, the location of diffusion inflection points, and the upper limit of diffusion. Accordingly, a number of studies on how to balance the dimensions of “efficient output” and “social equity and security” have pointed out that with the rapid development of technology, the depth and breadth of data applications continue to increase, which will involve changes in the economy, business, public administration, national security, scientific research, consumer rights and many other areas (Lazer et al., 2014). Specifically, as the degree of big data sharing increases, it increases consumer surplus and consumer utility, but at the same time, as data is co-generated by users in a data-driven model of firm competition, users buy the firm’s services and contribute data, and the firm invests in data infrastructure and collection.
Data collection improves services and benefits users, but users also face security issues such as privacy breaches (Cong and Mayer, 2023), which in turn incur privacy costs derived from data oversharing (Cong et al., 2021; Jones and Tonetti, 2020). With regard to the use and regulation of big data, e-commerce online platforms may use consumer information materialized by data elements to alter the “gloss” of product quality, causing consumers to misjudge the true quality of the product and inducing unwanted consumption behaviors (Acemoglu et al., 2025), Moreover, platforms will not only use the pricing mechanism of “big data to kill maturity” to make profits, but may also redistribute information in the form of “filter bubbles” for heterogeneous personal data in order to maximize platform engagement (Acemoglu et al., 2024). A study by Ding et al. (2024) argues that the stimulus effect on consumption growth of consumer benefits translated from data elements, such as the issuance of digital consumption vouchers or shopping subsidies, stems from the increase in consumer spending in the targeted category and not from crowding out of other categories of consumption expenditure. Madsen and Vellodi (2025) point out that a complete ban on data use may foster or inhibit innovation, depending on the demand for new products, and therefore more flexible policies that control when and what data are made available are needed to increase the effectiveness of data regulation. Some studies have incorporated information uncertainty into the “efficiency-equity” research framework, pointing out that information frictions connect data generation and economic activity, mainly in the following ways: micro-level uncertainty creates resource mismatches in the macro-system, making macro total factor productivity endogenous to micro agents’ data collection behavior (Farboodi and Veldkamp, 2021; David et al., 2016; Benhabib et al., 2016). This may be due to the widespread use of facilitated mobile communication tools and the availability of data information that allows for more random and rapid changes in individual behavior (Fabregas et al., 2025), which leads to the inability of traditional data selection mechanisms to accurately identify imperfect information and decision-making errors, as Gans (2025b) points out that in domains rich in dataization variety, AI’s analytical and decision-making capabilities performs well, while it is less credible in judgment-intensive data environments. Therefore, Caplin (2025) introduces new forms of data to identify agents’ preferences, beliefs, etc. by constructing data engineering models. Notably, Jones (2025) and Jones (2024) explore the potential for big data-enabled AI technologies to pose a threat to human existence while promoting economic growth at the level of heterogeneous agent utility acquisition.
To summarize, there are two mainstream conclusive studies in academia: the first one views data as a form of medium for information dissemination, which determines the source and value of information (Farboodi and Veldkamp, 2021); and the second one views data as an input factor, which determines technological advancement and output (Jones and Tonetti, 2020). And our data economy [^2] The framework of our paper is based on the perspective of “data economy”, which is different from digital economy. Data economy is a form of economic research centered on data as a factor of production, which centers on value creation through the production, distribution, exchange and consumption of data resources. Data economy emphasizes that data, as a non-rivalry, reusable resource, can be embedded in the production function to optimize the efficiency of resource allocation and enhance productivity. Its theoretical roots can be traced back to the breakthrough of the assumption of exogenous technology in neoclassical growth theory and the extension of knowledge spillovers in Rome’s (1990) endogenous growth theory. Digital economy refers to the new economic system reconstructed on the basis of Digital Technologies, covering the two dimensions of digital infrastructure, digital industrialization and industrial digitization. Its substance is the use of digital general-purpose technology, triggered by the total factor productivity increase and organizational model change, manifested in the expansion of the production boundary and the reduction of transaction costs. analysis refers to Jones and Tonetti (2020), viewing data as a factor of production to study its impact on the economy.
Marginal contributions to the article: First, we analyze the impact of dataization and technological progress on the dynamic general equilibrium of “consumption-capital” and summarize the conclusions about the development of data economy through numerical simulation and phase diagram analysis, filling the gap in the basic theory of this field. Secondly, most of the existing articles focus on the impact of dataization shock or the proliferation of data technologies on the economy, ignoring the role of dataization as a catalyst for technological progress from a theoretical perspective. Therefore, the article specifically explores the relationship between dataization and technological progress and argues for it using empirical analyses, providing recommendations for achieving faster, more persistent and more effective technological progress in the data economy.
2 The Model
2.1 The Production Function of Data
The Prerequisite. We assume that all consumers and firms in the macro-economy are homogeneous, i.e., representative consumers and representative firms. In this macro-economy, the consumption of representative consumers and the investment of representative firms will generate big data, and these data elements can be collected and applied to the study of consumer consumption behavior and firm investment behavior, and to the formation of algorithms and techniques. The total output of the macro-economy comes from the economic information generated by the decisions of representative consumers and firms, and as to what proportion of the total output can be legally dataized and applied to research, this paper sets out to show that this is an exogenous variable that varies with changes in digital economy policies, which will be determined by the policies set by the government. That is, policies that do not allow some of the data to be applied to algorithms for studying consumer and business behavior due to the dual factors of enhancing consumer welfare and protecting individual privacy, or, policies that stipulate that only no more than a certain percentage of the total output can be applied to the development of technology and innovation.
Assuming the Cobb-Douglas production function without data is given by the following equation:
$y(t)=(z(t)k(t))^{\alpha}l(t)^{\beta}$ (1)
where $y(t)$ is output, $z(t)$ is the technological attachment of capital, $k(t)$ is capital, and $l(t)$ is labor. $\alpha$ and $\beta$ are the output elasticities of capital and labor, respectively, with $\alpha\in(0,1)$ and $\beta\in(0,1)$.
Now, taking data as an element of the production function, based on Jones and Tonetti (2020), assume an exponential relationship $\eta$ between data $d(t)$ and technology $z(t)$. Introducing the exogenous variable $\theta$, assume that the policy stipulates that only $\theta$ proportions of output can be dataized and put into production as data elements, i.e.:
$\begin{gathered}z(t)=d(t)^{\eta},\eta\in(0,1)\\ d(t)=\theta y(t),\theta\in(0,1)\end{gathered}$
Solving by associating $y(t)$,$z(t)$,$d(t)$, we get:
$y(t)=(\theta^{\eta}k(t))^{\frac{\alpha}{1-\alpha\eta}}l(t)^{\frac{\beta}{1-\alpha\eta}}=\theta^{\frac{\alpha\eta}{1-\alpha\eta}}k(t)^{\frac{\alpha}{1-\alpha\eta}}l(t)^{\frac{\beta}{1-\alpha\eta}}$ (2)
It is important to note that the exponential transformation rate relationship between data and technology stems from the nonrivalry and cumulative effects of data. As a factor of production, data can be shared indefinitely without loss, and its value grows exponentially with the scope of its use. In the technical production function, the amount of data directly affects the production efficiency of technology through the exponential form:
$z(t)=d(t)^{\eta}$
$\eta=\left(\frac{logz(t)}{logd(t)}\right)$
Hence, $\eta$ reflects the importance of data to technology. As the amount of data increases, the output efficiency of technology increases at an exponential rate, forming a positive cycle: more data→higher technical efficiency→more output→generate more data. For example, the training of autonomous driving techniques requires a large amount of driving data. Assuming that an algorithm uses 1 million kilometers of driving data with 80% accuracy, when the amount of data is increased to 10 million kilometers (10 times), the accuracy may be increased to 95% due to the cumulative effect of non-competitive data (non-linear growth). This improvement is not linearly cumulative, but rather the data is trained by the algorithm to form an exponential optimization, reflecting the exponential transformation relationship between data and technology.
2.2 The Firm Problem
Endogenous Interest Rate. First consider the accounting profit of the firm and then consider the investment based on the accounting profit of the firm. Accounting profit is expressed as:
$\pi(k,w)=\max_{l}\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}k^{\frac{\alpha}{1-\alpha\eta}}l^{\frac{\beta}{1-\alpha\eta}}-wl\right\}$ (3)
FOC:
$\begin{gathered}l^{*}=\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}k^{\frac{\alpha}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\\ \pi(k,w)=\theta^{\frac{\alpha\eta}{1-\alpha\eta}}k^{\frac{\alpha}{1-\alpha\eta}}(l^{*})^{\frac{\beta}{1-\alpha\eta}}-wl^{*}\\ =\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}-w\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}k^{-\frac{\alpha}{\beta+\alpha\eta-1}}\\ =\underbrace{\pi(w)k^{\frac{\alpha}{1-\beta-\alpha\eta}}}_{\text{{accounting profit}}}\end{gathered}$
Which:
$\pi(w)=\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}-w\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}$ (4)
Further, the economic profit of the firm can be written as:
$\pi(r,w)=\max_{k}\left\{\pi(w)k^{\frac{\alpha}{1-\beta-\alpha\eta}}-rk\right\}$
Accordingly, the endogenous interest rate can be obtained:
$r=\frac{\alpha}{1-\beta-\alpha\eta}\pi(w)k^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}$ (5)
And:
$\pi(w)=\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}-w\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}$
Maximizing Capital Stock. The investment problem of the firm is as follows:
[…]
s.t.:
$\dot{k}(t)=i(t)-\delta k(t)$
$\pi(w(t))=\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w(t)\right]^{\frac{\beta}{\beta+\alpha\eta-1)}}-w(t)\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w(t)\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}$
The current-value Hamiltonian is written as:
$H=\underbrace{\pi(w(t))k(t)^{\frac{\alpha}{1-\beta-\alpha\eta}}}_{\text{{accounting profit}}}-\underbrace{i(t)}_{investment}-\underbrace{\frac{1}{2}a\left[\frac{i(t)}{k(t)}-\delta\right]^{2}k(t)}_{\text{{adjustment cost}}}+\underbrace{q(t)}_{multiplier}\underbrace{(i(t)-\delta k(t))}_{\dot{k}(t)}$ (7)
FOC:
$\begin{gathered}\frac{\partial H}{\partial i}=0\\ 1+a\left[\frac{i(t)}{k(t)}-\delta\right]=q(t)\\ \frac{i(t)}{k(t)}=\delta+\frac{1}{a}[q(t)-1]\\ \frac{\partial H}{\partial k}=\pi(w(t))\left(\frac{\alpha}{1-\beta-\alpha\eta}\right)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}+\frac{1}{2}a\left[\left(\frac{i(t)}{k(t)}\right)^{2}-\delta^{2}\right]-q(t)\delta\\ \dot{q}(t)=\frac{dq(t)}{dt}=(r+\delta)q(t)-\pi(w(t))\left(\frac{\alpha}{1-\beta-\alpha\eta}\right)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}-\frac{1}{2}a\left[\left(\frac{i(t)}{k(t)}\right)^{2}-\delta^{2}\right]\end{gathered}$
Given:
$\left(\frac{i(t)}{k(t)}\right)^{2}=\delta^{2}+\frac{1}{a^{2}}[q(t)-1]^{2}+\frac{2\delta}{a}[q(t)-1]$
Hence:
$\dot{q}(t)=(r+\delta)q(t)-\pi(w(t))\left(\frac{\alpha}{1-\beta-\alpha\eta}\right)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}-\frac{1}{2}a\left\{\frac{1}{a^{2}}[q(t)-1]^{2}+\frac{2\delta}{a}[q(t)-1]\right\}$
The first-order condition for investment can be written as:
$\underbrace{\dot{q}(t)=0}_{\text{{steady-state investment}}}$
Under steady-state conditions:
$k=\left\{\frac{(r+\delta)q-\frac{1}{2}a\left\{\frac{1}{a^{2}}[q(t)-1]^{2}+\frac{2\delta}{\alpha}[q(t)-1]\right\}}{\pi(w)\left(\frac{\alpha}{1-\beta-\alpha\eta}\right)}\right\}^{\frac{1-\beta-\alpha\eta}{\alpha+\beta+\alpha\eta-1}}$ (8)
Plugging $i(t)=\left\{\delta+\frac{1}{a}[q(t)-1]\right\}k(t)$:
$\begin{gathered}\dot{k}(t)=\left\{\delta+\frac{1}{a}[q(t)-1]\right\}k(t)-\delta k(t)\\ \frac{\dot{k}(t)}{k(t)}=\delta+\frac{1}{a}[q(t)-1]-\delta=\frac{1}{a}[q(t)-1]\end{gathered}$
When the capital stock is in steady state, $\dot{k}(t)=0$, $q=1$.
2.3 The Household Problem
Maximizing Utility.Assuming that the consumer utility function is a CRRA utility function, consumption is $c(t)$, the risk aversion coefficient $\sigma>1$, and the utility discount factor is $\rho$, the utility maximization problem is as follows:
[…]
s.t.:
$\begin{gathered}\dot{k}(t)=r(t)k(t)+w(t)l(t)-c(t)-\delta k(t)\\ r(t)=\frac{\alpha}{1-\beta-\alpha\eta}\pi(w)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}\\ c(t)\geq 0\end{gathered}$
According to equation $(4)$, $r(t)$ is the endogenous interest rate and comes from the firm’s investment problem. For the household, the current-value Hamiltonian is written as:
$H(k(t),c(t),\lambda(t))=\frac{c(t)^{1-\sigma}}{1-\sigma}+\lambda(t)[r(t)k(t)+w(t)l(t)-c(t)-\delta k(t)]$ (10)
FOC:
$\begin{gathered}\frac{\partial H}{\partial c}=0\\ c(t)^{-\sigma}=\lambda(t)\\ \dot{\lambda}(t)=\rho\lambda(t)-\frac{\partial H}{\partial k}\\ \frac{\dot{\lambda}(t)}{\lambda(t)}=\rho-r(t)+\delta\\ \frac{\dot{c}(t)}{c(t)}=\frac{r(t)-\rho-\delta}{\sigma}\end{gathered}$
Plugging $r(t)$, we get:
$\frac{\dot{c}(t)}{c(t)}=\frac{\frac{\alpha}{1-\beta-\alpha\eta}\pi(w)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}-\rho-\delta}{\sigma}$ (11)
Dynamic General Equilibrium Analysis. According to equation $(3)$, we get $l(t)=l^{*}$, and combining the formula $(3),(4),(5),(9),(11)$, We can get the dynamic general equilibrium system of our data economy:
$\begin{gathered}\frac{\dot{c}(t)}{c(t)}=\frac{\frac{\alpha}{1-\beta-\alpha\eta}\pi(w)k(t)^{\frac{\alpha+\beta+\alpha\eta-1}{1-\beta-\alpha\eta}}-\rho-\delta}{\sigma}\\ \frac{\dot{k}(t)}{k(t)}=\frac{\theta^{\frac{\alpha\eta}{1-\beta-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}k(t)^{\frac{\alpha}{1-\beta-\alpha\eta}}-c(t)-\delta k(t)}{k(t)}\\ \pi(w)=\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}-w\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}\end{gathered}$
When the system is in steady state $c^{*}$, $k^{*}$, $\dot{c}(t)=\dot{k}(t)=0$, Thus we can get analytic solutions for equilibrium consumption $c^{*}$ and equilibrium capital $k^{*}$:
$\begin{gathered}k^{*}=\left({\frac{(\rho+\delta)(1-\beta-\alpha\eta)}{\alpha\left\{\theta^{\frac{\alpha\eta}{1-\alpha\eta}}\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}-w\left[\left(\frac{1-\alpha\eta}{\beta}\right)\theta^{\frac{\alpha\eta}{\alpha\eta-1}}w\right]^{\frac{1-\alpha\eta}{\beta+\alpha\eta-1}}\right\}}}\right)^{\frac{1-\beta-\alpha\eta}{\alpha+\beta+\alpha\eta-1}}\\ c^{*}=\theta^{\frac{\alpha\eta}{1-\beta-\alpha\eta}}\left[(\frac{1-\alpha\eta}{\beta})w\right]^{\frac{\beta}{(\beta+\alpha\eta-1)}}(k^{*})^{\frac{\alpha}{1-\beta-\alpha\eta}}-\delta k^{*}\end{gathered}$
In the general equilibrium analysis of this paper, given $r=r(t)$, which is the endogenous interest rate given by the firm’s profit problem, the consumer solves the utility maximization problem and the firm solves the investment maximization problem according to $r(t)$.
3 Numerical Simulation and Phase Diagram Analysis
Parameter Setting.According to the analytical solution $(c^{*},k^{*})$,the parameters of this paper are set as follows: Assuming $0<\alpha+\beta\leq 1$, referring to the study of Jones and Tonetti (2020), the output elasticity of data factors in the production function is distributed between $0.03-0.12$. Therefore, we set the capital output elasticity $\alpha=0.6$ and the labor output elasticity $\beta=0.2$; $w$ is a parameter measuring the wage level, $w=1$; $\delta$ is the capital depreciation rate, $\delta=0.08$; and $\rho$ is the utility discount rate, $\rho=0.07$. And $\theta\in(0,1),\eta\in(0,1)$.
The three-dimensional simulation results are shown in Figures 1:
Figure 1: Three-dimensional illustration of the effects of $\theta$ and $\eta$ on $k^{*}$ and $c^{*}$.
From Figure 1-A: The effect of changes in $\eta$ and $\theta$ on equilibrium investment $k^{*}$ is linear. An increase in the data technology transformation rate $\eta$ promotes investment, while an increase in the proportion of total output dataization $\theta$ depresses investment. Thus, the boosting effect of $\eta$ on the equilibrium capital stock will gradually slow down with an increase in $\theta$. Figure 1-B shows that the impact of changes in the data technology conversion rate $\eta$ and the proportion of total output dataization $\theta$ on the equilibrium consumption $c^{*}$ is non-linear and has a “mixing” effect. $c^{*}$ is affected by $\eta$ at an inflection point, and the inflection point moves backward with increasing $\theta$. This implies that the promotion of consumption by $\eta$ is not persistent, and there is a possibility that the equilibrium consumption level will decrease at higher stages of data technology conversion. We interpret this as follows:
On the one hand, it is clear from Figure 1-A that technological progress can increase the marginal productivity of capital factors thereby driving the equilibrium capital stock, which is consistent with the rules of reality. And when the dataization ratio of total output increases, from an economic perspective, higher $\theta$ implies that more data are produced, and technology significantly improves and raises total factor productivity, at this time, firms are more biased toward data-driven rather than capital factor inputs for technological improvements, and technological improvement contributes to output partially substituting for capital, coupled with a shift in resources from physical capital investment to data production, leading to a decrease in capital required in the steady state. And, in an environment where technology is highly sensitive to data, i.e., the higher the level of $\eta$, the more pronounced is the impact of the equilibrium capital stock by the increase in the proportion of dataization. As can be seen from Figure 1-B, when $\eta$ is small, an increase in $\theta$ does not have a significant effect on equilibrium consumption. In contrast, when $\eta$ is large, an increase in $\theta$ has the effect of promoting equilibrium consumption. Since $\eta$ is the data-technology conversion rate, which represents the maturity of the mechanism for converting data into technology, the larger $\eta$ is, the more mature the conversion mechanism is, the more the same amount of data will be converted into a higher level of technological achievements, thus more efficiently enabling the transmission of initial data from the production side to the demand side of the individual. Consumers also utilize data-forming technologies to assist in decision-making and effective consumption. Therefore, with a mature mechanism of data technology conversion, increasing the proportion of total macroeconomic output dataization can stimulate consumption growth.
On the other hand, Figure 1-B shows that the overall impact of $\eta$ on equilibrium consumption shows an inverted “U”-shaped trend, which implies that transforming big data into technology is not always effective in stimulating consumption growth, i.e., there exists a threshold for the impact of an increase in the data-technology conversion rate on equilibrium consumption, and the size of this threshold is related to The size of this threshold is related to the proportion of total output dataization $\theta$. Specifically, when $\theta$ is larger, in the domain of definition of $\eta$, the range of promoting effect of increasing $\eta$ on $c^{*}$ will be more persistent, and the threshold of the inverted “U”-shaped trend will be more backward, and gradually evolve into “S”-shaped logistic curve. This may be due to the scale of data is the basis for the establishment of the data-technology transformation mechanism (Learning from Big Data). Therefore, the proportion of total output dataization increases, which means that after the total output is dataized, the volume of data that can be used for technological transformation also increases, and the space and resources for “learning” from data will be expanded, the more reliable the data-technology transformation mechanism is, the higher the success rate of data transformation into technology, and the more accurate the prediction of technology for consumer behavior, thus driving the “resilience” of effective consumption growth.
Further examining Figure 1, we also find that both for $k^{*}$ and $c^{*}$, when $\eta$ is at a higher level ($0.6<\eta<1)$, there is a positive correlation between $\theta$ and $\eta$ at the same level of equilibrium (The Same Color Area in the Figure 1). Therefore, when consumption or capital stock is at the same level of steady state, increasing the macroeconomic total output dataization ratio $\theta$ can effectively promote technical progress $\eta$. Specifically as shown in Figure 2. This will be the theoretical basis for our empirical analysis.
Figure 2: Positive correlation between $\theta$ and $\eta$ at the same level of equilibrium
Phase Diagram Analysis. Figures 3, 4 further present the dynamic general equilibrium subject to shocks from changes in $\eta$ and $\theta$ in the macro-economy of this paper, using the phase diagram approach:
Figure 3: Shocks to dynamic general equilibrium from increased $\eta$
Figure 3 presents the impact of an increase in the data-technology conversion rate $\eta$ on consumption and capital stock in general dynamic equilibrium for a constant dataization ratio $\theta$ of total macroeconomic output: The black curve represents the case before the positive shock to $\eta$, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The blue curve represents the case at the beginning of the positive shock to $\eta$, with a small increase in $\eta$, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The red curve represents the case at the middle to late stage of the positive shock to $\eta$, with a larger increase in $\eta$, $E_{3}$ represents the equilibrium point, and $P_{3}$ represents the inflection point in that case; The three inflection points $P_{1}$,$P_{2}$,$P_{3}$ are covariant since the aggregate output dataization ratio $\theta$ is constant. When $\eta$ increases, equilibrium capital stock will increase linearly, and the phase diagram is shifted to the right. The blue phase diagram indicates that an increase in $\eta$ before the inflection point value will drive equilibrium consumption growth, and the red phase diagram indicates that an increase in $\eta$ after the inflection point value will slightly reduce equilibrium consumption.
Figure 4: Shocks to dynamic general equilibrium from increased $\theta$
As can be seen from Figures 1-B, the impact of increasing dataization ratio $\theta$ on equilibrium consumption is not apparent in the stage of lower $\eta$. Therefore, the impact of increasing macroeconomic total output dataization ratio $\theta$ on consumption and capital stock in the dynamic general equilibrium in Figure 4 is analyzed for the stage of higher $\eta$: The black curve represents the case before the positive shock of theta, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The red curve represents the case when $\eta$ is higher as $\theta$ increases, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The two inflection points, $P_{1}$,$P_{2}$, increase sequentially to the right as the dataization ratio of total output $\theta$ continues to increase. As $\theta$ increases, the equilibrium capital stock will be linearly lower, and the phase diagram is shifted left. The red phase diagram indicates that expanding the proportion of dataization at the mature stage of data-technology conversion will be able to effectively stimulate consumption growth.
From figure 3 we can know that when the dataization ratio $\theta$ is fixed, it means that the threshold of $\eta$ to promote $c^{*}$ growth is fixed, and at this time a small increase in $\eta$ will effectively promote $c^{*}$ growth, while once the threshold is exceeded, continuing to raise $\eta$ will not be conducive to $c^{*}$ growth. And regardless of the magnitude of $\eta$ raising, its effect on $k^{*}$ is always promoting. From Figure 4 we can know that when eta is at a higher stage, raising the proportion of dataization $\theta$ will be able to effectively promote the $c^{*}$ growth and reduce $k^{*}$, at the same time, the threshold value of the effect of $\eta$ raising on the growth of $c^{*}$ will also increase because of $\theta$ increasing.
4 Empirical Evidence and Analysis
From the numerical simulation analysis and the results of Figure 2, it can be seen that an increase in the proportion of total output dataization $\theta$ can promote technological progress, i.e., an increase in $\eta$, when the equilibrium capital stock $k^{*}$ and the equilibrium consumption $c^{*}$ are at the steady-state level, i.e., at fixed values. Accordingly, we present the core proposition of the empirical analysis section:
Proposition: Open data policies can drive digital technological advances and contribute to the development of the digital economy.
We design a simple and effective DID experiment to prove the empirical proposition. Macroeconomic data of China from 2000 to 2022 are selected: First, the digital economy development index (Digital) of 216 prefecture-level cities and four municipalities in China is taken as one of the explanatory variables, in which Digital is measured in terms of Internet development and digital financial inclusion, taking into account the availability of relevant data at the city level. Internet penetration rate, related employees, related output and cell phone penetration rate are adopted as the measurement indicators of Internet development, and the actual contents corresponding to the above four indicators are: the number of Internet broadband access users per 100 people, the proportion of employees in the computer service and software industry in the employees in urban units, the total amount of telecommunication services per capita, and the number of cell phone subscribers per 100 people. The raw data for the above indicators were obtained from the China Urban Statistical Yearbook. The “China Digital Inclusive Finance Index” is used as a measure of inclusive finance, which is jointly compiled by the Digital Finance Research Center of Peking University and Ant Financial Services Group. Finally, the Digital variable will be calculated by standardizing and down scaling the data of the above five indicators through Principal Component Analysis to obtain the Comprehensive Development Index of the Digital Economy; Second, the corresponding 220 cities’ “digital economy-related utility model patents granted” (Tech) (unit:10,000 pieces) from 2000 to 2022 is used as one of the explanatory variables, representing an indicator of digital technological progress, and the patent data are from the State Intellectual Property Office.
This paper uses the Staggered Difference-in-Difference model to explore how Digital and Tech change after the implementation of open data policies. Using the open data pilot cities as the core experimental subjects, the product of the policy implementation time dummy variable and the experimental group dummy variable constitutes the DID. The time dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 for this year and the following years, otherwise it will be 0. The experimental group dummy variable is: if a provincial region has conducted an open data pilot in a certain year, the region will be assigned a value of 1 from this year onward, and prior to this year the region is assigned a value of 0, which is the control group. In this paper, two macro policies related to data openness are selected, the first one is “Establishment of Regional Government Data Open Platform” (DID1), and the second one is “Establishment of Regional Comprehensive Pilot Zones for Cross-border E-commerce” (DID2). The time cut-off points for the implementation of both policies are taken from public documents on the official websites of the central and local governments.
The prerequisite for our experimental core proposition to hold is the need for capital stock and consumption in steady state, so total regional fixed asset investment/10,000 yuan (Capital) and total retail sales of consumer goods/10,000 yuan (Consumption) are chosen as control variables. In addition, We selected gross domestic product/10,000 yuan (GDP), regional household population/10,000 people (People), number of industrial enterprises above scale/1 enterprise (Company), total wages of employees on board/10,000 yuan (Salary), total freight volume/10,000 tons (Freight), total telecommunication services/10,000 yuan (Teleservice), unemployment insurance participants/1 people (Insurance) as the control variables [^3] All control variables were treated by taking common logarithms.. All the data comes from the Official Website of Regional Government and the China Urban Statistical Yearbook [^4] The number of study samples matched to the two types of policies will be different due to limitations in the timing of policy implementation and the availability of sample data: The time interval for the DID1 correspondence sample is 2000-2022, with the earliest implementation date in 2012 and the latest in 2021. The time interval for DID2 is 2010-2022, with the earliest implementation date in 2016 and the latest implementation date in 2022. And the current period of policy implementation will be deleted from all DID experiments..
The results of the benchmark experiment [^5] Standard errors in parentheses: ${}^{*}p<0.05$,${}^{**}p<0.01$,${}^{***}p<0.001$ are shown in Table 1, and the results of the policy effects experiment are shown in Figure 5, Figure 6, Figure 7, Figure 8, they show the shocks of dataization policy on technical progress variables under two-way fixed effects. From them, we find that dataization policies can significantly promote technological progress, consistent with the core proposition of the article:
Columns (1)-(3) show the impact of the first type of dataization policy, i.e., “government data platform establishment”, on technological progress: it is found that the first type of dataization policy has a significant effect on the development of the digital economy and the number of patents granted in the digital economy. Although the promotion effect of the first type of policy on the development of digital economy is no longer significant after adding the time fixed effect, we can know through Figure 5 that this is because there is a lag in the promotion effect of the first type of policy on the development of digital economy, which may be related to the implementation of the policy of data openness, the establishment of databases and data management platforms need time. After the 4th period of the first type of policy shock, the effects of such policies on the development of digital economy are all significantly positive. Figure 6 presents the impact effect of the first type of policy on the number of patents granted in the digital economy. The coefficients are not significant before the policy shock, and after the policy shock, the coefficients are all significantly positive, which fully demonstrates that the first type of dataization policy can effectively promote the enhancement of patent granting volume of digital economy.
Columns (4)-(6) show the impact of the second type of dataization policy, i.e., the “establishment of comprehensive cross-border e-commerce pilot zones”, on technological progress: from which it is found that the second type of dataization policy has, in the case of two-way fixed effects, a significant boosting effect. As can be seen from Figure 7, the promotion effect of the second type of policy on digital economy development is similar to the case of the first type of policy, with a lag period, and it is only after the 4th period of the second type of policy shock that it is able to significantly promote the development of the digital economy, and it is therefore reasonable to assume that this is due to the time needed for the establishment of databases and data management platforms. As can be seen from Figure 8, the second type of policy has a very good positive shock effect on the number of patents granted in the digital economy, so it is reasonable to believe that the second type of policy can significantly increase the number of patents granted in the digital economy.
Table 1: Benchmark Regression Results
Dependent Variable
(1) (2) (3) (4) (5) (6)
Independent Variable Digital Digital Tech Digital Digital Tech
DID1 $0.021^{***}$ $0.001$ $0.034^{**}$
$(0.002)$ $(0.002)$ $(0.013)$
DID2 $0.018^{***}$ $0.005^{*}$ $0.151^{***}$
$(0.002)$ $(0.002)$ $(0.015)$
Capital $-0.003^{***}$ $-0.010^{***}$ $-0.269^{***}$ $-0.007^{***}$ $-0.005$ $-0.232^{***}$
$(0.000)$ $(0.002)$ $(0.014)$ $(0.000)$ $(0.003)$ $(0.020)$
Consumption $0.024^{***}$ $-0.001$ $0.045^{***}$ $0.007^{**}$ $-0.004$ $0.040^{*}$
$(0.002)$ $(0.002)$ $(0.011)$ $(0.003)$ $(0.003)$ $(0.020)$
GDP $0.011^{***}$ $0.002$ $0.021^{**}$ $0.006^{***}$ $-0.001$ $0.002$
$(0.001)$ $(0.001)$ $(0.008)$ $(0.001)$ $(0.001)$ $(0.009)$
Company $0.019^{***}$ $-0.001$ $0.097^{***}$ $0.028^{***}$ $0.009^{*}$ $0.176^{***}$
$(0.003)$ $(0.002)$ $(0.016)$ $(0.005)$ $(0.004)$ $(0.033)$
Population $-0.018^{***}$ $0.012^{***}$ $0.208^{***}$ $-0.005^{*}$ $0.001$ $0.100^{***}$
$(0.001)$ $(0.003)$ $(0.018)$ $(0.002)$ $(0.004)$ $(0.026)$
Salary $0.003^{***}$ $0.003^{***}$ $0.014^{***}$ $-0.000$ $0.001$ $0.010^{*}$
$(0.001)$ $(0.001)$ $(0.004)$ $(0.001)$ $(0.001)$ $(0.004)$
Freight $-0.007^{***}$ $-0.005^{*}$ $-0.063^{***}$ $-0.007^{***}$ $-0.002$ $-0.077^{***}$
$(0.000)$ $(0.002)$ $(0.013)$ $(0.000)$ $(0.003)$ $(0.019)$
Teleservice $0.008^{***}$ $0.001$ $-0.008$ $0.002^{*}$ $0.001$ $0.001$
$(0.000)$ $(0.001)$ $(0.005)$ $(0.001)$ $(0.001)$ $(0.007)$
Insurance $0.009^{***}$ $0.009^{***}$ $0.169^{***}$ $0.005^{***}$ $0.001$ $0.058^{***}$
$(0.000)$ $(0.001)$ $(0.009)$ $(0.000)$ $(0.002)$ $(0.013)$
Constant $0.107^{***}$ $0.189^{***}$ $-0.103^{*}$ $0.216^{***}$ $0.294^{***}$ $0.226^{*}$
$(0.006)$ $(0.006)$ $(0.041)$ $(0.014)$ $(0.014)$ $(0.102)$
City $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Year $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
$R^{2}$ 0.817 0.880 0.369 0.657 0.731 0.291
Number 4922 4922 4953 2792 2792 2813
Figure 5: Shock from DID1 to Digital
Figure 6: Shock from DID1 to Tech
Figure 7: Shock from DID2 to Digital
Figure 8: Shock from DID2 to Tech
5 Conclusion
In summary, the present paper identifies three analysis theories for the data economy, we attempt to complement and refine the fundamental economic theory of productivity growth, technological progress, and dynamic general equilibrium in the field of data economics :
Firstly, technological advances have the potential to significantly promote the equilibrium capital stock of firms. That is, the increase of $\eta$ can lead to an increase in $k^{*}$. However, the degree of data openness has been demonstrated to be inversely correlated with equilibrium capital stock, that is, the increase of $\theta$ will lead to a decrease in $k^{*}$, which means the greater the degree of data openness, the less the equilibrium capital stock, and the stronger the substitution effect of data on capital factors. Secondly, for the equilibrium consumption of households, the impact of technological progress on equilibrium consumption is a typical U-shaped curve, and the policy regarding data openness will make it more persistent, gradually transforming it from U-shaped curve to S-shaped curve. That means the increase of data openness $\theta$ could assist in enhancing the threshold of technological progress $\eta$ for equilibrium consumption $c^{*}$ growth. Finally, we generalize the nonlinear effects of data openness policy on the impact of technological progress on dynamic general equilibrium to an important theory of data economy: data openness could profoundly promote the progress of digital technology and the development of digital economy. The proposition is formulated through numerical simulations, and four DID experiments are designed to demonstrate the core ideas. The empirical results from our experiments are consistent and significant with the theoretical analysis.